OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! Find the specific solution of the differential equation dy dx equals the quotient of 2 times y and x squared with condition y(−2) = e. a. y= -1-2/x b. y= e^-2/x c. y= -e^1/x d. none of these

zepdrix (zepdrix):

(2y)/x^2 i think

OpenStudy (aryana_maria2323):

\[\frac{ dy }{ dx }= \frac{ 2y }{ x^2 }\]

OpenStudy (across):

Just separate variables, integrate, and exponentiate to get rid of the natural logarithm.

OpenStudy (aryana_maria2323):

What do you mean? Can you show me in steps? I have to show my work and I have no idea where to start. I have two other problems like this so once I learn how to do this one I can do the other ones.

OpenStudy (across):

Find the general solution: \[ \begin{align} \frac{dy}{dx}=\frac{2y}{x^2}&\implies\frac{dy}{y}=\frac{2}{x^2}dx\\ &\implies\ln y=-\frac{2}{x}+c\\ &\implies y=ce^{-\frac{2}{x}} \end{align} \] Apply the initial value: \[ e=ce^{-\frac{2}{-2}}=e\implies c=1. \] Express your final solution: \[ y=e^{-\frac{2}{x}}. \]

OpenStudy (across):

Typo under the initial value part, but you get the idea.

OpenStudy (aryana_maria2323):

Okay I understand up until the ln. Where did ln come from?

OpenStudy (across):

\[\int\frac1x\,dx=\ln x+c\]

OpenStudy (aryana_maria2323):

Ohh okay I see now.

OpenStudy (aryana_maria2323):

Can you help me work through the other one?

OpenStudy (across):

post it

OpenStudy (across):

Make a new question. I won't go over it here.

OpenStudy (aryana_maria2323):

Okay and thank you.